# Reassessing aspects of the photon’s LQG-modified dispersion relations

### v1

@@ I Introduction
Nowadays, there is a broad diversity of phenomena where quantum gravity effects could be observable, among which there is an increasing interest in the category of the so-called tabletop experiments for quantum gravity. In this paper, we focus on the phenomenon of the energy-dependent time of arrival of cosmic photons and neutrinos from distant sources [7, 8]. LQG is a relevant candidate to explain how the speed of light in vacuum may display an energy-dependent behavior. This happens because, at the Planck energy scale, one of the basic principles of General Relativity (GR), namely, Lorentz invariance, is broken by virtue of the strong vacuum fluctuations. This implies the formulation of field equations in the electromagnetic sector that incorporates the effects of LQG [7]. This yields different approaches and the development of various formulations [9, 10, 8]. Let us recall that Gamma-Ray Bursts (GRBs) [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] and Active Galactic Nuclei (AGNs) are examples of extremely energetic phenomena and appear as an abundant source of highly energetic particles, like cosmic photons and neutrinos. These ultra-energetic particles may reveal Lorentz-invariance violation (LIV) as an observable effect. LQG appears consequently as a viable candidate to provide a theoretical basis to explain this sort of phenomenon. GRBs constitute a remarkable category of astrophysical structure to allow the detection of quantum-gravitational effects. The travel time differences accumulated after photons propagate from long distances are a key element to unravel these effects . Because it is hard to distinguish two massless particles with different energies that travel from distant regions where GRBs occur, it becomes necessary to observe short and intense bursts traveling large distances to guarantee a good experimental time resolution.
@@
The focus of our present paper is however the vacuum dispersion relation emerging from the LQG-corrected set of Maxwell equations. We are concerned with a high-energy model, so that the photons exhibit astrophysical-scale contributions [22, 23]. This opens up the possibility of analyzing deviations to be compared with standard dispersion relations, potentially indicating imprints of quantum-gravitational effects. Other fundamental quantities that could serve as probes for analyzing Quantum Gravity (QG) include photons exhibiting a vacuum speed different from the value predicted by General Relativity, or refractive indices deviating from the case of usual (Maxwellian) light. All the modifications of the equations lead to alterations in known events and optical properties of light in vacuum , such as birefringence, which is the dependence of the refractive index on the polarization, dichroism and group-velocuty dispersion, among others. These facts re-inforce the possibility of detecting QG effects that emerge in high-energy events.
With the new class of modified photon dispersion relations we shall be assessing, we examine how they modify the expression for the wavelength shift of the photon in a Compton scattering. Originally, this effect results from extensive investigations of the scattering of X-rays by matter [24]. This process played a crucial role to the synthesis of Quantum Mechanics and Electrodynamics, with culminated with the complete quantitative explanation proposed by Dirac, with his formulation of Relativistic Quantum Mechanics [25]. While the Compton effect highlights quantum interactions at the particle level, LQG extends these principles to spacetime, striving to describe the behavior of gravity at the most fundamental scales. Understanding such connections deepens our comprehension of the Universe, bridging quantum phenomena with the fabric of spacetime. The results to be presented later on shall explore the consequences of this union, potentially unveiling deeper structures of spacetime that integrate the quantum behavior of particles with the gravitational dynamics of the universe.
The outline of our paper is as follows. In Sec. II, we briefly describe the introduction of the electromagnetic field in a scenario of LQG to set up the general presentation of the model wee shall be working with. Next, in Section III, we study classical aspects of the system under consideration. In some situations, we work in the approximation of linearized magnetic field, for in one of those modified Maxwell equations there appears a nonlinear term in the magnetic field. Section IV focus on the Compton effect reassessed with the LQG-modified dispersion relations for the photon. In this case, we switch off the nonlinear term in the magnetic field and follow a path to identify the LQG effects in the expression for the shift of the photon wavelengths. To conclude, in Sec. V, we present our Final Comments and General Discussions.
@@ II A quick glance at the model
In order to place our discussion into a clear context, we start off by providing a summary of the model under consideration, governed by the Hamiltonian density given below:
@@
[…]
We specify some of the quantities present in Eq. (II). $Q^{2}$ is the electromagnetic coupling constant, $l_{p}\approx 1.6\times 10^{-35}m$ is the Planck length. The characteristic length $\mathcal{L}$ is constrained by the relation $l_{P}\ll\mathcal{L}\leq\lambda$, where $\lambda$ is the de Broglie wavelength. The other characteristic length, $\mathcal{L}$, has a maximum value at $\mathcal{L}=k^{-1}$. This makes clear that we are actually working within an effective theory. Finally, $\Upsilon$ may depend on the helicity of the particle under consideration [27, 28, 29] and $\theta_{i}$’s are dimensionless parameters of order one or they are extremely small, close to zero [7, 30]; $a,b$ are spatial tensor indices. To make a clear notation, we eliminate all underlines in the electromagnetic parameters and, from Eq. (II), we rewrite the vectors in bold. The field equations are given as shown below:
@@ III Electromagnetic Radiation
We consider the dimensionless parameters $\bar{\theta}_{i}$ obey the conditions below:
@@
Similarly to what we normally do in classical electromagnetism, we can take the cross product of equations Eqs. (II) and (5) with the fields E and B, respectively. By proceeding as indicated, we can readily arrive at LQG-corrected expressions for the energy density and the Poynting vector:
@@
Eqs. (10) and (14) reduce respectively to the usual Poynting vector and energy density of the electromagnetic field, whenever we remove the LQG parameters. The continuity equation of the Poynting vector is kept unchanged, but the radiated energy, as shown above, has news terms coming from LQG: each $\bar{\theta}_{i}$ contributes independent with one term to the the spatial form, as we can see in the terms $\textbf{S}_{1},\textbf{S}_{2}\ \mbox{and}\ \textbf{S}_{3}$. To our sense, at this point, it is important to remark that the nonlinear term in the magnetic field is the only term modifies the expression for the electromagnetic energy density. From the Poynting vector given in Eq. (9), by taking its time derivative, we can work out a continuity equation which allows us to write down the purely space components of the energy-momentum tensor, the so-called normal stress and shear stress :
@@
$-2i\bar{\theta}_{4}\bar{\theta}_{8}(\textbf{k}\times\bm{\zeta})_{i}\cdot b_{0j}\textbf{k}^{2}.$ (29)
@@
[…]
@@
The $\psi=-\bar{\theta}_{4}[A_{\gamma}-\bar{\theta}_{3}(\textbf{k})^{2}]\cdot|\textbf{k}|^{2}\cdot|\bm{\zeta}|^{2}\sin^{2}{\varphi}$ and $\pm$ signs correspond to the different polarizations of the photon. If the $\bm{\zeta}$ or the linear terms $\bar{\theta}_{3}=\bar{\theta}_{8}=\bar{\theta}_{7}=0$, we obtain the same results given in the work of Ref. [7]. The group velocity can also be directly derived from the dispersion relation. Differentiating the latter yields the expression for the group velocity:
@@
±$\pm\frac{1}{4}\Biggl{(}4k^{4}\Biggl{\{}k^{2}\left([A_{\gamma}-\bar{\theta}_{3}(\textbf{k})^{2}]-\bar{\theta}_{4}\frac{\zeta^{2}}{2}\right)^{2}+\left([A_{\gamma}-\bar{\theta}_{3}(\textbf{k})^{2}]-\frac{\zeta^{2}}{2}\right)\cdot\bar{\theta}_{4}\cdot[\bm{\zeta}\cdot(\textbf{k}\times(\textbf{k}\times\bm{\zeta}))]\Biggl{\}}\bar{\theta}_{8}^{2}+\psi^{2}\Biggl{)}^{-1}\cdot$
@@
$+\left([A_{\gamma}-\bar{\theta}_{3}(\textbf{k})^{2}]\!-\!\frac{\zeta^{2}}{2}\right)\bar{\theta}_{4}[\bm{\zeta}\cdot(\textbf{k}\times(\textbf{k}\times\bm{\zeta}))]\biggl{\}}\bar{\theta}_{8}^{2}\!+\!\psi^{2}\Biggl{)}\Biggl{)}^{-1/2}.$ (33)
@@ IV LQG corrections to the kinematics of the Compton effect
In this Section, we shall address to the corrections to the Compton effect stemming from the LQG-corrected photon dispersion relations. Before we initiate the computation, it is a fundamental point to highlight that we do not use the magnetic field splitting given by Eq. (7) in this Section. For our purposes in the present Section, we shall instead work with Eqs. (II) and (5), where we will be disregarding the nonlinear term in the magnetic field. Thus, based on the Eqs. (2) and (4) that describe the source-free situation, again we consider plane wave solutions:
$\textbf{E}=\textbf{e}_{0}e^{i(\textbf{k}\cdot\textbf{x}-\omega t)},\hskip 8.5359pt\textbf{B}=\textbf{b}_{0}e^{i(\textbf{k}\cdot\textbf{x}-\omega t)},\hskip 8.5359ptk=|\textbf{k}|.$ (34)
We obtain a electromagnetic field
@@
$(\textbf{k}\times\textbf{B}_{0})(A_{\gamma}-\bar{\theta}_{3}\textbf{k}^{2})-i\bar{\theta}_{8}\textbf{k}^{2}\cdot\textbf{B}_{0}+w\textbf{E}_{0}=0$ (36)
[…]
$\omega=c\textbf{k}\left[A_{\gamma}+\theta_{3}(\textbf{k}\ell_{p})^{2}\pm\theta_{8}(\textbf{k}\ell_{p})\right].$ (37)
The equation above is the same as the one coming from Eq. (III), whenever $\bm{\zeta}=0$. Using the relation above in the derivation of the Compton effect (for a better understanding of the procedure, see [35] and [34]), through the energy and momentum conservation, we find that the difference in wavelengths between the scattered and incident photons, taking into account LQG effects, is given by:
@@
Now, let us simplify the equation above. First, let us disregard the third term on the right-hand side, as it is of second order in the scale $\ell_{p}$. We are interested in keeping only first-order term in the LQG parameters. We adopt the moving scale, which is related to the characteristic scale $\mathcal{L}$ and the momentum by $\mathcal{L}=k^{-1}$. Our result for the shift in the photon wavelength is cast below:
$\lambda^{\prime}-\lambda=$ $\lambda_{c}(1-\cos\theta)-\theta_{7}\lambda_{c}\Biggl{\{}\biggl{[}\frac{\lambda^{\prime}}{\lambda}(\ell_{p}k)^{2+2\Upsilon}+\frac{\lambda}{\lambda^{\prime}}(\ell_{p}k^{\prime})^{2+2\Upsilon}\biggr{]}-\biggl{[}(\ell_{p}k)^{2+2\Upsilon}(\ell_{p}k^{\prime})^{2+2\Upsilon}\biggl{]}$
$+\frac{1}{\lambda_{c}}\left[\lambda^{\prime}\left(\ell_{p}k\right)^{2+2\Upsilon}+\lambda(\ell_{p}k^{\prime})^{2+2\Upsilon}\right]\Biggr{\}}\pm 4\pi\theta_{8}\ell_{p}\lambda_{c}\left[\left(\frac{1}{\lambda}+\frac{1}{\lambda^{\prime}}\right)+\left(\frac{\lambda^{\prime}}{\lambda^{2}}+\frac{\lambda}{\lambda^{\prime 2}}\right)+\frac{1}{\lambda_{c}}\left(\frac{\lambda^{\prime}}{\lambda}+\frac{\lambda}{\lambda^{\prime}}\right)\right].$ (39)
Notice that at the right side of the Eq. (IV) we have the helicity term, where we choose the value $\Upsilon=-1/2$, in such a way that our framework can be qualitatively reproduced. Thus, considering that $k=2\pi/\lambda$, the equation becomes:
$\lambda^{\prime}-\lambda=$ $\lambda_{c}(1-\cos\theta)-2\pi\theta_{7}\ell_{p}\lambda_{c}\biggl{[}\left(\frac{\lambda^{\prime}}{\lambda^{2}}+\frac{\lambda}{\lambda^{\prime 2}}\right)-\left(\frac{1}{\lambda}+\frac{1}{\lambda^{\prime}}\right)+\frac{1}{\lambda_{c}}\left(\frac{\lambda^{\prime}}{\lambda}-\frac{\lambda}{\lambda^{\prime}}\right)\biggr{]}$ (40)
±$\pm 4\pi\theta_{8}\ell_{p}\lambda_{c}\left[\left(\frac{1}{\lambda}+\frac{1}{\lambda^{\prime}}\right)+\left(\frac{\lambda^{\prime}}{\lambda^{2}}+\frac{\lambda}{\lambda^{\prime 2}}\right)+\frac{1}{\lambda_{c}}\left(\frac{\lambda^{\prime}}{\lambda}+\frac{\lambda}{\lambda^{\prime}}\right)\right].$
@@
The best constraint on the rotation angle $\theta_{8}$ in the literature [40] can be transformed into a restriction $\theta_{8}\leq 10^{-16}$, approximately, for which we can simply set $\theta_{8}$ to zero, such that $\ell_{P}k\leq 10^{-13}$, approximately, and therefore, $\theta_{8}\ell_{P}k\sim(\ell_{P}k)^{2}$. Consequently, the Compton effect with the LQG corrections becomes
$\lambda^{\prime}-\lambda=\lambda_{c}(1-\cos\theta)-2\pi\theta_{7}\ell_{p}\lambda_{c}\biggl{[}\left(\frac{\lambda^{\prime}}{\lambda^{2}}+\frac{\lambda}{\lambda^{\prime 2}}\right)-\left(\frac{1}{\lambda}+\frac{1}{\lambda^{\prime}}\right)+\frac{1}{\lambda_{c}}\left(\frac{\lambda^{\prime}}{\lambda}-\frac{\lambda}{\lambda^{\prime}}\right)\biggr{]}.$ (42)

### v2

@@ I Introduction
Nowadays, there is a broad diversity of phenomena where quantum gravity effects could be observable, among which there is an increasing interest in the category of the so-called tabletop experiments for quantum gravity. In this paper, we focus on the phenomenon of the energy-dependent time of arrival of cosmic photons and neutrinos from distant sources [7, 8]. LQG is a relevant candidate to explain how the speed of light in vacuum may display an energy-dependent behavior. This happens because, at the Planck energy scale, one of the basic principles of General Relativity (GR), namely, Lorentz invariance, is broken by virtue of the strong vacuum fluctuations. This implies the formulation of field equations in the electromagnetic sector that incorporates the effects of LQG [7]. This yields different approaches and the development of various formulations [9, 10, 8]. Let us recall that Gamma-Ray Bursts (GRBs) [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] and Active Galactic Nuclei (AGNs) are examples of extremely energetic phenomena and appear as abundant source of highly energetic particles, like cosmic photons and neutrinos. These ultra-energetic particles may reveal Lorentz-invariance violation (LIV) as an observable effect. LQG appears consequently as a viable candidate to provide a theoretical basis to explain this sort of phenomenon. GRBs constitute a remarkable category of astrophysical structure to allow the detection of quantum-gravitational effects. The travel time differences accumulated after photons propagate from long distances are a key elements to unravel these effects. Because it is hard to distinguish two massless particles with different energies that travel from distant regions where GRBs occur, it becomes necessary to observe short and intense bursts traveling large distances to guarantee a good experimental time resolution.
@@
The focus of our present paper is however the vacuum dispersion relation emerging from the LQG-corrected set of Maxwell equations. We are concerned with a high-energy model, so that the photons exhibit astrophysical-scale contributions [22, 23]. This opens up the possibility of analyzing deviations to be compared with standard dispersion relations, potentially indicating imprints of quantum-gravitational effects. Other fundamental quantities that could serve as probes for analyzing Quantum Gravity (QG) include photons exhibiting a vacuum speed different from the value predicted by General Relativity, or refractive indices deviating from the case of usual (Maxwellian) light. All the modifications of the equations lead to alterations in known events and optical properties of light in vacuum , such as birefringence, which is the dependence of the refractive index on the polarization, dichroism and group-velocity dispersion, among others. These facts re-inforce the possibility of detecting QG effects that emerge in high-energy events.
With the new class of modified photon dispersion relations we shall be assessing, we examine how they modify the expression for the wavelength shift of the photon in a Compton scattering. Originally, this effect results from extensive investigations of the scattering of X-rays by matter [24]. This process played a crucial role in the synthesis of Quantum Mechanics and Electrodynamics, which culminated in the complete quantitative explanation proposed by Dirac through his formulation of Relativistic Quantum Mechanics [25]. While the Compton effect highlights quantum interactions at the particle level, LQG extends these principles to spacetime, striving to describe the behavior of gravity at the most fundamental scales. Understanding such connections deepens our comprehension of the Universe, bridging quantum phenomena with the fabric of spacetime. The results to be presented later on shall explore the consequences of this union, potentially unveiling deeper structures of spacetime that integrate the quantum behavior of particles with the gravitational dynamics of the universe.
The outline of our paper is as follows. In Sec. II, we briefly describe the introduction of the electromagnetic field in a scenario of LQG to set up the general presentation of the model we shall be working with. Next, in Section III, we study classical aspects of the system under consideration. In some situations, we work in the approximation of linearized magnetic field, for in one of those modified Maxwell equations there appears a nonlinear term in the magnetic field. Section IV focus on the Compton effect reassessed with the LQG-modified dispersion relations for the photon. In this case, we switch off the nonlinear term in the magnetic field and follow a path to identify the LQG effects in the expression for the shift of the photon wavelengths. To conclude, in Sec. V, we present our Final Comments and General Discussions.
@@ II A quick glance at the model
In order to place our discussion into a clear context, we start off by providing a summary of the model under consideration. The Hamiltonian formulation of Loop Quantum Gravity (LQG) effects on electromagnetic theory was explored in the work of Ref. [7] is given below:
@@
$\;+$ $\theta_{3}l_{P}^{2}(\underline{B}^{a}\nabla^{2}\underline{B}_{a}+\underline{E}^{a}\nabla^{2}\underline{E}_{a})+\theta_{2}l_{P}^{2}\underline{E}^{a}\partial_{a}\partial_{b}\underline{E}^{b}$
$\;+$ $\theta_{8}l_{P}[\underline{\vec{B}}\cdot(\nabla\times\underline{\vec{B}})+\underline{\vec{E}}\cdot(\nabla\times\underline{\vec{E}})]$
@@
We specify some of the quantities present in Eq. (II). $Q^{2}$ is the electromagnetic coupling constant, $l_{p}\approx 1.6\times 10^{-35}m$ is the Planck length. The characteristic length $\mathcal{L}$ is constrained by the relation $l_{P}\ll\mathcal{L}\leq\lambda$, where $\lambda$ is the de Broglie wavelength. The characteristic length, $\mathcal{L}$, has a maximum value at $\mathcal{L}=k^{-1}$. This makes clear that we are actually working within an effective theory. Finally, $\Upsilon$ may depend on the helicity of the particle under consideration [27, 28, 29] and $\theta_{i}$’s are dimensionless parameters of order one or they are extremely small, close to zero [7, 30]; $a,b$ are spatial tensor indices. To make a clear notation, we eliminate all underlines in the electromagnetic parameters and, from Eq. (II), we rewrite the vectors in bold. The field equations are given as shown below:
@@ III Electromagnetic Radiation
We consider the dimensional parameters $\bar{\theta}_{i}$ obey the conditions below:
@@
This may introduce a simplification, because each parameter, $\bar{\theta}_{i}$, can be associated with a different physical phenomenon. Similarly to what we normally do in classical electromagnetism, we can take the cross product of equations Eqs. (II) and (5) with the fields E and B, respectively. By proceeding as indicated, we can readily arrive at LQG-corrected expressions for the energy density and the Poynting vector:
@@
Eqs. (10) and (14) reduce respectively to the usual Poynting vector and energy density of the electromagnetic field, whenever we remove the LQG parameters. The continuity equation of the Poynting vector is kept unchanged, but the radiated energy, as shown above, has new terms coming from LQG: each $\bar{\theta}_{i}$ contributes independent with one term to the the spatial form, as we can see in the terms $\textbf{S}_{1},\textbf{S}_{2}\ \mbox{and}\ \textbf{S}_{3}$. To our sense, at this point, it is important to remark that the nonlinear term in the magnetic field is the only term modifies the expression for the electromagnetic energy density. From the Poynting vector given in Eq. (9), by taking its time derivative, we can work out a continuity equation which allows us to write down the purely space components of the energy-momentum tensor, the so-called normal stress and shear stress :
@@
$-2i\bar{\theta}_{4}\bar{\theta}_{8}(\textbf{k}\times\bm{\zeta})_{i}\cdot\zeta_{j}\textbf{k}^{2}.$ (28)
@@
whose determinant is given by $\mbox{det M}=\alpha^{3}+c^{2}(\textbf{u}\cdot\textbf{v})\cdot(\gamma\textbf{s}\cdot\textbf{v}+\beta(\textbf{u}\cdot\textbf{v}))+\alpha^{2}\beta\textbf{u}^{2}+\alpha c(c\textbf{v}^{2}+\gamma\textbf{s}\cdot(\textbf{v}\times(\textbf{v}\times\textbf{s})))$. Now, we can satisfy the condition $\mbox{det M}=0$ what possibility to find the modified dispersion relation
@@
The term $\psi=-\bar{\theta}_{4}[A_{\gamma}-\bar{\theta}_{3}(\textbf{k})^{2}]\cdot|\textbf{k}|^{2}\cdot|\bm{\zeta}|^{2}\sin^{2}{\varphi}$ is a scalar, but it manifestss the anisotropy of the model, in that photons present propagation speeds that are no longer constant at the Planck scale. This is precisely due to the dependence that the $\psi$ parameter has on $\sin\varphi$, where this is the angle between the vectors $\textbf{k}\cdot\bm{\zeta}$. The $\pm$ sign in the dispersion relation is an indication of the expected phenomenon of vacuum birefringence. If the $\bm{\zeta}$ or the linear terms $\bar{\theta}_{3}=\bar{\theta}_{8}=\bar{\theta}_{7}=0$, we obtain the same results given in the work of Ref. [7]. The group velocity can also be directly derived from the dispersion relation. Differentiating the latter expression yields the expression for the group velocity:
@@
$\pm\frac{1}{4}\Biggl(4k^{4}\Biggl\{k^{2}\left([A_{\gamma}-\bar{\theta}_{3}(\textbf{k})^{2}]-\bar{\theta}_{4}\frac{\zeta^{2}}{2}\right)^{2}+\left([A_{\gamma}-\bar{\theta}_{3}(\textbf{k})^{2}]-\frac{\zeta^{2}}{2}\right)\cdot\bar{\theta}_{4}\cdot[\bm{\zeta}\cdot(\textbf{k}\times(\textbf{k}\times\bm{\zeta}))]\Biggl\}\bar{\theta}_{8}^{2}+\psi^{2}\Biggl)^{-1}\cdot$
@@
$+\left([A_{\gamma}-\bar{\theta}_{3}(\textbf{k})^{2}]\!-\!\frac{\zeta^{2}}{2}\right)\bar{\theta}_{4}[\bm{\zeta}\cdot(\textbf{k}\times(\textbf{k}\times\bm{\zeta}))]\biggl\}\bar{\theta}_{8}^{2}\!+\!\psi^{2}\Biggl)^{1/2}\Biggl)^{-1/2}.$ (32)
@@ IV LQG corrections to the kinematics of the Compton effect
In this Section, we shall address to the corrections to the Compton effect stemming from the LQG-corrected photon dispersion relations. Before we initiate the computation, it is a fundamental point to highlight that we do not use the magnetic field splitting given by Eq. (7) in this Section. For our purposes in the present Section, we shall instead work with Eqs. (II) and (5), where disregard the nonlinear term in the magnetic field. By then considering plane wave solutions for the electric and magnetic fields,
$\textbf{E}=\textbf{e}_{0}e^{i(\textbf{k}\cdot\textbf{x}-\omega t)},\hskip 8.5359pt\textbf{B}=\textbf{b}_{0}e^{i(\textbf{k}\cdot\textbf{x}-\omega t)},\hskip 8.5359ptk=|\textbf{k}|,$ (33)
we get:
@@
$(\textbf{k}\times\textbf{B}_{0})(A_{\gamma}-\bar{\theta}_{3}\textbf{k}^{2})-i\bar{\theta}_{8}\textbf{k}^{2}\cdot\textbf{B}_{0}+w\textbf{E}_{0}=0.$ (35)
Notice that we do not consider the nonlinear contribution in the magnetic field of (5), so that Eqs. (25) and (III)are different from the Eqs. (33) and (34) previously obtained. From Eq. (27) and the condition $\mbox{det \ M}=0$, it is possible to obtain the modified dispersion relation; it takes the form:
$\omega=c|\textbf{k}|\left[A_{\gamma}+\theta_{3}(|\textbf{k}|\ell_{p})^{2}\pm\theta_{8}(|\textbf{k}|\ell_{p})\right].$ (36)
The equation above is the same as the one coming from Eq. (III), whenever $\bm{\zeta}=0$. We adopt this dispersion relation in the derivation of the kinematics of the Compton effect (for a better understanding of the procedure, see [35] and [34]. Through the energy and momentum conservation, we find that the difference in wavelengths between the scattered and incident photons, taking into account LQG effects, is given by:
@@
Now, let us simplify the equation above. First, let us disregard the third term on the right-hand side, as it is of second order in the scale $\ell_{p}$. We are interested in keeping only first-order terms in the LQG parameters. We adopt the moving scale, which is related to the characteristic scale $\mathcal{L}$ and the momentum by $\mathcal{L}=k^{-1}$. Our result for the shift in the photon wavelength is cast below:
$\lambda^{\prime}-\lambda=$ $\lambda_{e}(1-\cos\theta)-\theta_{7}\lambda_{e}\Biggl\{\biggl[\frac{\lambda^{\prime}}{\lambda}(\ell_{p}k)^{2+2\Upsilon}+\frac{\lambda}{\lambda^{\prime}}(\ell_{p}k^{\prime})^{2+2\Upsilon}\biggr]-\biggl[(\ell_{p}k)^{2+2\Upsilon}(\ell_{p}k^{\prime})^{2+2\Upsilon}\biggl]$
$+\frac{1}{\lambda_{e}}\left[\lambda^{\prime}\left(\ell_{p}k\right)^{2+2\Upsilon}+\lambda(\ell_{p}k^{\prime})^{2+2\Upsilon}\right]\Biggr\}\pm 4\pi\theta_{8}\ell_{p}\lambda_{e}\left[\left(\frac{1}{\lambda}+\frac{1}{\lambda^{\prime}}\right)+\left(\frac{\lambda^{\prime}}{\lambda^{2}}+\frac{\lambda}{\lambda^{\prime 2}}\right)+\frac{1}{\lambda_{e}}\left(\frac{\lambda^{\prime}}{\lambda}+\frac{\lambda}{\lambda^{\prime}}\right)\right].$ (38)
Here, we define the Compton wavelength $\lambda_{e}=h/mc$, where $m$ is the electron mass. Notice that, at the right side of the Eq. (IV), we have the helicity term, where we choose the value $\Upsilon=-1/2$, in such a way that our framework can be qualitatively reproduced. Thus, considering that $k=2\pi/\lambda$, the equation becomes:
$\lambda^{\prime}-\lambda=$ $\lambda_{e}(1-\cos\theta)-2\pi\theta_{7}\ell_{p}\lambda_{e}\biggl[\left(\frac{\lambda^{\prime}}{\lambda^{2}}+\frac{\lambda}{\lambda^{\prime 2}}\right)-\left(\frac{1}{\lambda}+\frac{1}{\lambda^{\prime}}\right)+\frac{1}{\lambda_{e}}\left(\frac{\lambda^{\prime}}{\lambda}-\frac{\lambda}{\lambda^{\prime}}\right)\biggr]$ (39)
$\pm 4\pi\theta_{8}\ell_{p}\lambda_{e}\left[\left(\frac{1}{\lambda}+\frac{1}{\lambda^{\prime}}\right)+\left(\frac{\lambda^{\prime}}{\lambda^{2}}+\frac{\lambda}{\lambda^{\prime 2}}\right)+\frac{1}{\lambda_{e}}\left(\frac{\lambda^{\prime}}{\lambda}+\frac{\lambda}{\lambda^{\prime}}\right)\right].$
@@
The best constraint on the rotation angle $\theta_{8}$ in the literature [40] can be transformed into a restriction $\theta_{8}\leq 10^{-16}$. Thus, the $\theta_{8}\ell_{p}\lambda_{e}$ component is of the order of $10^{-63}$. So, we also cancel out the term multiplied by these components. Consequently, the expression for the Compton effect with the LQG corrections becomes as follows:
$\lambda^{\prime}-\lambda=\lambda_{e}(1-\cos\theta)-2\pi\theta_{7}\ell_{p}\lambda_{e}\biggl[\left(\frac{\lambda^{\prime}}{\lambda^{2}}+\frac{\lambda}{\lambda^{\prime 2}}\right)-\left(\frac{1}{\lambda}+\frac{1}{\lambda^{\prime}}\right)+\frac{1}{\lambda_{e}}\left(\frac{\lambda^{\prime}}{\lambda}-\frac{\lambda}{\lambda^{\prime}}\right)\biggr].$ (41)
@@
To extract an estimate of the LQG contribution to the kinematics of the Compton effect, let us notice that, from Eq. (41), the shift on the wavelength of the photon, $\Delta\lambda=\lambda^{\prime}-\lambda$, splits into two pieces. If we denote by $\Delta_{c}\lambda$ the usual Compton shift, whereas $\Delta_{LQG}$ stands for the effect of LQG on $\Delta\lambda$,
$\Delta\lambda=\Delta_{c}\lambda+\Delta_{LQG}\lambda,$ (42)
we may evaluate how they compete by deriving the ratio $\Delta_{LQG}\lambda/\Delta_{c}\lambda$.
According to the inspection carried out by Li and Ma in the work of Ref. [36], the factor $\theta_{7}\ell_{p}$, in natural units, turns out to be negative and estimated as
$\theta_{7}\ell_{p}\sim-2.8\times 10^{-18}GeV^{-1};$ (43)
in length units this corresponds to
$\theta_{7}\ell_{p}\sim-5.49\times 10^{-32}cm.$ (44)
Now, by replacing $\lambda^{\prime}$ as $\lambda^{\prime}=\lambda+\Delta\lambda$ in Eq. (41) and by considering that $\Delta_{LQG}\lambda\ll\lambda$ (terms in higher powers are neglected), yields:
$\frac{\Delta_{LQG}\lambda}{\Delta_{c}\lambda}\simeq-4\pi\frac{\theta_{7}\ell_{p}}{\lambda}.$ (45)
We highlight that this result does not depend on the Compton wavelength of the particle that scatters the radiation. Let us also notice that the LQG contribution is more sensitive to the hight frequencies/short wavelengths. For typical X-ray radiations,
$\Delta_{LQG}\lambda\Big|_{X}\sim 10^{-21}\Delta_{c}\lambda\Big|_{X};$ (46)
for radiation in the deep gamma spectrum $(\nu\sim 10^{25}\mbox{Hz})$, the estimate is
$\Delta_{LQG}\lambda\Big|_{\gamma}\sim 10^{-14}\Delta_{c}\lambda\Big|_{\gamma}.$ (47)